The refractive index of glass is 1.5. Find the speed of light in glass.

Grade:12th pass

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
We have,

\mu = \frac{speed \, of \, light \, in\, vaccum}{speed \, of \, light \, in \, the\, material}

Thus, speed of light in glass,

= \frac{speed\, of\, light\, in\, vaccum\, }{\mu }

\frac{3.0\times 10^{8}ms^{-1}}{1.5} = 2.0\times 10^{8} ms^{-1}